OpenStudy (anonymous):

find the second derivative of x^2-y^2=36 in terms of x and y

OpenStudy (polaris_s0i):

if you want in terms of x and y you want to do implicit differentiation: \[x^2 - y^2 = 36\] \[2xdx - 2ydy = 0\] \[2dx^2 - 2dy^2 = 0\]

OpenStudy (anonymous):

yea i know dy/dx=x/y but idk how to take the second derivative of it

OpenStudy (polaris_s0i):

this is the total derivative, given by: \[f^\prime=\frac{\partial}{\partial{x}}+\frac{\partial}{\partial{y}}\] \[f^{\prime\prime}=\frac{\partial^2}{\partial{x}}+2\frac{\partial^2}{\partial{x}\partial{y}}+\frac{\partial^2}{\partial{y}}\]

OpenStudy (polaris_s0i):

I am assuming that is what they want?

OpenStudy (anonymous):

they want the second derivative

OpenStudy (polaris_s0i):

if they want d^2y that is different, this assumes both variables are independent.

OpenStudy (anonymous):

yes thats what they want

OpenStudy (polaris_s0i):

so it should just be \[2dx^2 - 2dy^2 = 0\]

OpenStudy (anonymous):

idk it says the answer is -36/y^3

OpenStudy (polaris_s0i):

ok, so they aren't independent then.. I was just thinking that they weren't cause they aren't wrting \[f(x,y) = x^2 - y^2\]

OpenStudy (anonymous):

yes like that

OpenStudy (polaris_s0i):

no x, that's odd... gonna hit the whiteboard a sec.

zepdrix (zepdrix):

\[\Large x^2+y^2=36\]\[\Large 2x+2yy'=0\quad\to\quad y'=\frac{x}{y}\]Then to find the second derivative we apply the quotient rule:\[\Large y''=\frac{y-xy'}{y^2}\]And then to simplify it down to the answer key I guess we need to plug some stuff in.. hmm Do you understand how I applied quotient rule there?

OpenStudy (anonymous):

no the equation is x2-y2=36

zepdrix (zepdrix):

Ahh sorry typo :( The first derivative should be correct though. I meant to write x^2-y^2.

zepdrix (zepdrix):

But did you understand the quotient rule process? :U I know it can be a little tricky.

OpenStudy (polaris_s0i):

ah I see yes

OpenStudy (anonymous):

yes i understand, so it will be \[\frac{ y-x(\frac{ x }{ y}) }{ y ^{2}}\]

zepdrix (zepdrix):

Looks good so far!

OpenStudy (anonymous):

idk what comes next

OpenStudy (polaris_s0i):

right so you get: \[\frac{y^2 - x^2}{y} * \frac{1}{y^2}\] \[\frac{y^2-x^2}{y^3}\] which is the negative of the original equation. \[\frac{-36}{y^3}\]

OpenStudy (anonymous):

how did you get y^2-x^2?

OpenStudy (polaris_s0i):

because: \[\frac{y - x(\frac{x}{y}))}{y^2}\] \[\frac{\frac{y^2}{y} - \frac{x^2}{y}}{y^2}\] \[\frac{\frac{y^2 - x^2}{y}}{y^2}\] \[\frac{y^2 - x^2}{y}*\frac{1}{y^2}\]

OpenStudy (anonymous):

i understand why x^2 but idk how you get y^2?

OpenStudy (polaris_s0i):

because I'm multiplying by y/y to get a common denominator

OpenStudy (anonymous):

oooh okay got it!

OpenStudy (polaris_s0i):

\[1 * y = \frac{y}{y}*\frac{y}{1} = \frac{y^2}{y}\]

OpenStudy (polaris_s0i):

cool, good question... I learned quite a bit as well :)

OpenStudy (anonymous):

thank you so much!

OpenStudy (polaris_s0i):

no problem, good @zepdrix stepped in though :)